OpenStudy (anonymous):

Find a power series representation for the rational function using differentiation or partial fractions. f(x) = x^8 / (9+x^9 )^2

OpenStudy (anonymous):

\[f(x)= \sum_{n=1}^{\infty} = ?\]

OpenStudy (anonymous):

or maybe both?

OpenStudy (anonymous):

they want the function represented as a power series with n starting at 1

OpenStudy (anonymous):

\[\frac{x^9}{(9+x^9)^2}=\frac{1}{9+x^9}-\frac{9}{(9+x^9)^2}\]

OpenStudy (anonymous):

that's what i started doing but apparently it might easier to begin with a simpler function like 1/ (9+x^9 )

OpenStudy (anonymous):

and derive that

OpenStudy (anonymous):

that one you can do like \(\frac{1}{1+x}\)

OpenStudy (anonymous):

oh i see! i think that would make it easier

OpenStudy (anonymous):

so for that one you'd get \[\sum_{n=0}^{\infty} (-1/9)^n x ^{9n}\]

OpenStudy (anonymous):

oooh hold the phone you have two different problems listed

OpenStudy (anonymous):

oh sorry! it's the original problem with x^8

OpenStudy (anonymous):

\[f(x) = \frac{ x ^{8} }{ (9+x^9)^2 }\] then integrate and get \[-\frac{1}{9}\frac{1}{9+x^9}\]

OpenStudy (anonymous):

write that one as a power series, then differentiate term by term

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (-1) (1/9)^{n+1}x ^{9n}\]

OpenStudy (anonymous):

i think that's right, but i'm not sure how to make it so that n=1

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i'm stuck on differentiation term by term

OpenStudy (anonymous):

\[\frac{1}{9+x^9}=\frac{1}{9}\frac{1}{1+\frac{x^9}{9}}\] \[=\frac{1}{9}\left (1-\frac{x}{9}+\frac{x^2}{9^2}-...\right)\]

OpenStudy (anonymous):

the sigma notation is too hard for me